This question is a word problem using the quadratic formula A rectangular poool has dimensions of 40 ft and 60 ft. If the pool has a patio area around it that equals the area of the pool, how wide is the distance from the pool edge to the patio edge.
Answer by ankor@dixie-net.com(22740)

This question is a word problem using the quadratic formula A rectangular pool has dimensions of 40 ft and 60 ft. If the pool has a patio area around it that equals the area of the pool, how wide is the distance from the pool edge to the patio edge.
:
Let x = the width of the patio around the pool
:
The dimension of the pool and the patio:
(2x+40) by (2x+60)
:
Area of the pool 40*60 = 2400 q/ft, Total area of both: 4800 sq/ft
:
Total Area equation:
(2x+40)(2x+60) = 4800
FOIL
4x^2 + 200x + 2400 = 4800
:
4x^2 + 200x + 2400 - 4800 = 0
:
4x^2 + 200x - 2400 = 0
:
Simplify, divide by 4;
x^2 + 50x - 600 = 0
:
Factor:
(x+60)(x-10) = 0
:
The positive solution:
x = 10 ft is the width of the patio
;
Check solution by finding the total area: (40+20) * (60+20) = 4800
:
You can use the quadratic formula if you wish: a=1, b=50, c=-600